OpenStudy (anonymous):

How many right triangles have a hypotenuse that measures 2x+3 meters and legs that measure 2x-5 meters and x+7 meters? What are the dimensions of the triangle(s)?

OpenStudy (aum):

Use Pythagoras Theorem: \((2x+3)^2 = (2x-5)^2 + (x+7)^2\) Expand, simplify. It will be a quadratic equation. Solve for x.

OpenStudy (anonymous):

how would I expand? Do I move (2x+3)^2 to the right side? @aum

OpenStudy (aum):

\((a+b)^2 = a^2 + 2ab + b^2\)

OpenStudy (aum):

\( (2x+3)^2 = (2x-5)^2 + (x+7)^2 \\ 4x^2 + 12x + 9 = 4x^2 - 20x + 25 + x^2 + 14x + 49 \\ 4x^2 + 12x + 9 = 5x^2 - 6x + 74 \\ 5x^2 - 6x + 74 - 4x^2 -12x - 9 = 0 \\ x^2 -18x +65 = 0 \\ (x-5)(x-13) = 0 \\ x = 5, ~~ x = 13 \) Put x = 5: hypotenuse = 2x + 3 = 2*5 + 3 = 13; leg 1 = 2x - 5 = 2*5 - 5 = 5; leg 2 = x + 7 = 5 + 7 = 12. So the sides are 13, 5, 12.

OpenStudy (aum):

Put x = 13: hypotenuse = 2x + 3 = 2*13 + 3 = 29; leg 1 = 2x - 5 = 2*13 - 5 = 21; leg 2 = x + 7 = 13 + 7 = 20. So the sides are 29, 21, 20.